# Weighted Chernoff information and optimal loss exponent in context-sensitive hypothesis testing

### 09 Mar 2026 v1

@@ 2.2 Weighted affinities and Chernoff information
$\rho^{\mathrm{w}}_{\alpha}(p,q):=\int_{\cal X}\varphi(x)\,p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}\mu(x),$ (2.6)
@@
$D^{\mathrm{w}}_{B,\alpha}(p,q):=-\ln\rho^{\mathrm{w}}_{\alpha}(p,q).$ (2.7)
@@
$D^{\mathrm{w}}_{C}({\mathbb{P}},{\mathbb{Q}})=\max_{\alpha\in[0,1]}\left[-\ln\int_{\cal X}\varphi(x)\,p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}\mu(x)\right]=\max_{\alpha\in[0,1]}D^{\mathrm{w}}_{B,\alpha}(p,q).$ (2.8)
A maximiser $\alpha^{*}=\alpha^{*}(p,q)$ in (2.8) is called the optimal Chernoff parameter.
@@
3 Main results
@@ 3.1 Asymptotics of the optimal sum of losses
and that Proposition 2.1 yields (2.3). The next theorem identifies its exact logarithmic asymptotic rate with the weighted Chernoff information from Definition 2.4.
@@
Consider the binary hypotheses $H_{0}:\ X_{1}^{n}\sim\mathbb{P}^{\otimes n}$ versus $H_{1}:\ X_{1}^{n}\sim\mathbb{Q}^{\otimes n}$, under Assumption 1.1. Let $\mathcal{D}$ be the class of (possibly randomised) decision rules $D:{\cal X}^{n}\to[0,1]$, and let $\alpha_{\varphi}(D)$, $\beta_{\varphi}(D)$ be defined by (2.1)–(2.2). Assume that $p,q>0$ $\mu$-a.e. and that $\rho^{\mathrm{w}}_{\alpha}(p,q)\in(0,\infty)$ for all $\alpha\in[0,1]$.
@@
$L_{n}^{*}=\exp\{-nD_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})+o(n)\}.$ (3.1)
@@
For any $\alpha\in[0,1]$, $\min\{a,b\}\leq a^{\alpha}b^{1-\alpha}$, hence by factorisation,
@@
Taking $\inf_{\alpha\in[0,1]}$ yields $\liminf_{n\to\infty}-\frac{1}{n}\ln L_{n}^{*}\geq D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})$.
Fix $\alpha^{*}\in\arg\min_{\alpha\in[0,1]}\rho^{\mathrm{w}}_{\alpha}(p,q)$ and set $r_{\alpha^{*}}(x):=\varphi(x)p(x)^{\alpha^{*}}q(x)^{1-\alpha^{*}}/\rho^{\mathrm{w}}_{\alpha^{*}}(p,q)$. Let $S_{n}=\sum\limits_{i=1}^{n}\ln\frac{p(X_{i})}{q(X_{i})}$. A change of measure gives
$\begin{array}[]{l}L_{n}^{*}=(\rho^{*}_{\alpha^{*}}(p,q))^{n}\mathbb{E}_{{r^{*}}_{\alpha^{*}}^{\otimes n}}\Big[e^{-\alpha^{*}S_{n}}\mathbf{1}(S_{n}<0)+e^{(1-\alpha^{*})S_{n}}\mathbf{1}(S_{n}\geq 0)\Big].\end{array}$ (3.2)
The bracket is bounded by $1$.
Let $F(\alpha):=\ln\rho^{\mathrm{w}}_{\alpha}(p,q)$. Under the regularity assumptions of [5], $m:=\mathbb{E}_{r_{\alpha^{*}}}[\ln(p(X)/q(X))]=F^{\prime}(\alpha^{*})$ exists, and since $\alpha^{*}$ minimises $F$ on $[0,1]$, we have $m=0$ if $\alpha^{*}\in(0,1)$, while $m\leq 0$ if $\alpha^{*}=0$ and $m\geq 0$ if $\alpha^{*}=1$. Hence $S_{n}/n\to m$ in $r_{\alpha^{*}}$-probability. For any $\varepsilon>0$, on the event $\{|S_{n}/n-m|\leq\varepsilon\}$ the bracket is $\geq e^{-\varepsilon n}$, while the event has probability 1o(1)1-o(1). Therefore the expectation in brackets equals $\exp\{o(n)\}$.
Hence $L_{n}^{*}=\exp\{-nD_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})+o(n)\}$. For a more detailed treatment under additional regularity conditions we refer to [5]. ∎
@@
${\rm TV}_{\varphi}(\mathbb{P}^{\otimes n},\mathbb{Q}^{\otimes n})=\frac{1}{2}\Big((E_{\varphi}(p))^{n}+(E_{\varphi}(q))^{n}\Big)-\exp\{-nD^{\mathrm{w}}_{C}(\mathbb{P},\mathbb{Q})+o(n)\},\penalty 10000\ \penalty 10000\ n\to\infty,$ (3.3)
@@
Combine the identity (2.5) with (3.1). ∎
@@
The weighted $\alpha$-skewed Bhattacharyya distance (2.7) appears in many papers, see e.g. [7]. Definition 2.4 shows that the weighted Chernoff information divergence is the maximally skewed weighted Bhattacharyya distance.
@@ 3.2 Exponential-family representation and uniqueness of $\alpha^{*}$
$Z_{pq}(\alpha):=\int_{\cal X}\varphi(x)\,p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}\mu(x)=\rho^{\mathrm{w}}_{\alpha}(p,q),$ (3.4)
@@
${\cal E}_{pq}=\left\{(pq)_{\alpha}(x):=\frac{\varphi(x)\,p(x)^{\alpha}q(x)^{1-\alpha}}{Z_{pq}(\alpha)}\;:\;\alpha\in[0,1]\right\}.$ (3.5)
@@
Set $t(x):=\ln\frac{p(x)}{q(x)}$ and $k_{pq}(x):=\ln\varphi(x)+\ln q(x)$. Then $(pq)_{\alpha}$ admits the exponential-family form
$(pq)_{\alpha}(x)$ $=\exp\!\left\{\alpha\,t(x)-F_{pq}(\alpha)+k_{pq}(x)\right\},$ (3.6)
$F_{pq}(\alpha):$ $=\ln Z_{pq}(\alpha)=-D^{\mathrm{w}}_{B,\alpha}(p,q).$ (3.7)
@@
The log-normaliser $F_{pq}$ is convex on $[0,1]$; if $\ln\frac{p}{q}$ is not $\mu$-a.e. constant on $\{\varphi>0\}$, then $F_{pq}$ is strictly convex and the maximiser $\alpha^{*}$ in (2.8) is unique. By Hölder’s inequality,
@@
Finally, note that $Z_{pq}(1)=E_{\varphi}(p)$ and $Z_{pq}(0)=E_{\varphi}(q)$, hence
@@
so ${\cal E}_{pq}$ is an exponential arc between the tilted versions of $\mathbb{P}$ and $\mathbb{Q}$. By the definition, the following identities hold:
$\begin{array}[]{l}D^{\mathrm{w}}_{C}(p,q)=D^{\mathrm{w}}_{B,\alpha^{*}(p,q)}(p,q)=D^{\mathrm{w}}_{B,\alpha^{*}(q,p)}(q,p)=D^{\mathrm{w}}_{C}(q,p).\end{array}$ (3.8)
@@ 3.3 Weighted Bregman divergence and information-geometric identities
Weighted KL divergence and weighted Bregman divergence.
@@
$p_{\theta}(x)=\exp\{\theta^{\mathrm{T}}t(x)-F(\theta)+k(x)\},\qquad x\in{\cal X}.$ (3.9)
For a density $r$ set $E_{\varphi}(r):=\int_{\cal X}\varphi(x)\,r(x)\,\mathrm{d}\mu(x)$ and write $E_{\varphi}(\theta):=E_{\varphi}(p_{\theta})$.
@@
$\hat{F}(\theta):=\ln\int_{\cal X}\varphi(x)\,e^{\theta^{\mathrm{T}}t(x)+k(x)}\,\mathrm{d}\mu(x)=F(\theta)+\ln E_{\varphi}(\theta).$ (3.10)
@@
$p_{\theta}^{*}(x)=\frac{\varphi(x)\,p_{\theta}(x)}{E_{\varphi}(\theta)}.$ (3.11)
@@
$D^{\mathrm{w}}_{\rm KL}(p\|q):=\int_{\cal X}\varphi(x)\,p(x)\,\ln\frac{p(x)}{q(x)}\,\mathrm{d}\mu(x),$ (3.12)
@@
$B^{\mathrm{w}}_{\varphi,F}(\theta_{1},\theta_{2}):$ $=e^{\hat{F}(\theta_{2})-F(\theta_{2})}\Big[F(\theta_{1})-F(\theta_{2})-(\theta_{1}-\theta_{2})^{\mathrm{T}}\nabla\hat{F}(\theta_{2})\Big]$ (3.13)
@@
Remark 3.6.
For exponential families, the weighted KL divergence admits the Bregman representation $D^{\mathrm{w}}_{\rm KL}(p_{\theta_{1}}\|p_{\theta_{2}})=B^{\mathrm{w}}_{\varphi,F}(\theta_{2},\theta_{1})$; see [5, Proposition 4.1].
@@
Assume that the integral in (3.12) is well-defined in $(-\infty,\infty]$. Then for any $\theta_{1},\theta_{2}\in\Theta$,
$D^{\mathrm{w}}_{\rm KL}(p_{\theta_{1}}\|p_{\theta_{2}})=B^{\mathrm{w}}_{\varphi,F}(\theta_{2},\theta_{1}).$ (3.14)
@@
This identity is stated in [5, Proposition 4.1]; we give a short derivation for completeness. By (3.9),
@@
Substituting this into (3.12) yields
@@
Hence $\int\varphi tp_{\theta_{1}}\,\mathrm{d}\mu=E_{\varphi}(\theta_{1})\nabla\hat{F}(\theta_{1})$, and therefore
@@
(a) Weighted one-parameter identity. Assume $d=1$ (one-parameter case) and let $\theta_{i}^{*}:=F^{\prime}(\theta_{i})$. Then the following weighted analogue of the classical Bregman duality holds:
$B^{\mathrm{w}}_{\varphi,F}(\theta_{1},\theta_{2})=B^{\mathrm{w}}_{\varphi,F^{*}}(\theta_{2}^{*},\theta_{1}^{*})-(\theta_{1}-\theta_{2})\,\ln E_{\varphi}(\theta_{2})+(\theta_{2}^{*}-\theta_{1}^{*})\,\ln E_{\varphi}(\theta_{1}),$ (3.15)
@@
$B_{F}(\theta_{0},\theta_{1})=F(\theta_{0})+F^{*}(\theta_{1}^{*})-\theta_{0}^{\mathrm{T}}\theta_{1}^{*},\qquad\theta_{1}^{*}=\nabla F(\theta_{1}),$ (3.16)
@@
Part (a) is a weighted extension of the classical duality $B_{F}(\theta_{1},\theta_{2})=B_{F^{*}}(\theta_{2}^{*},\theta_{1}^{*})$ and follows by combining the weighted representation (3.13) with Legendre relations; see also [5]. Part (b) is standard. ∎
Weighted Chernoff/Bhattacharyya quantities inside an exponential family.

### 09 May 2026 v3

@@ 2.2 Weighted Affinities and Chernoff Information
$\rho^{\mathrm{w}}_{\alpha}(p,q):=\int_{\cal X}\varphi(x)\,p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}\mu(x),$ (11)
@@
$D^{\mathrm{w}}_{B,\alpha}(p,q):=-\ln\rho^{\mathrm{w}}_{\alpha}(p,q).$ (12)
@@
$D^{\mathrm{w}}_{C}({\mathbb{P}},{\mathbb{Q}})=\max_{\alpha\in[0,1]}\left[-\ln\int_{\cal X}\varphi(x)\,p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}\mu(x)\right]=\max_{\alpha\in[0,1]}D^{\mathrm{w}}_{B,\alpha}(p,q).$ (13)
A maximiser $\alpha^{*}=\alpha^{*}(p,q)$ in (13) is called the optimal Chernoff parameter.
@@
3 Asymptotics and Information-Geometric Identities
Before stating the main theorem, we separate the two optimisations that appear throughout this section and should not be conflated. The total loss $L_{n}(D)=\alpha_{\varphi}(D)+\beta_{\varphi}(D)$ is a non-negative functional of the decision rule and is minimised over $D\in\mathcal{D}$, giving the optimal loss $L_{n}^{*}=\inf_{D\in\mathcal{D}}L_{n}(D)$. The map $\alpha\mapsto-\ln\rho^{\mathrm{w}}_{\alpha}(p,q)$ is a non-negative concave functional of the skewing parameter and is maximised over $\alpha\in[0,1]$, giving the weighted Chernoff information $D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})=\sup_{\alpha\in[0,1]}[-\ln\rho^{\mathrm{w}}_{\alpha}(p,q)]$. Theorem 3.1 below connects the two via the single-letter asymptotic $L_{n}^{*}=\exp\{-nD_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})+o(n)\}$.
@@ 3.1 Asymptotics of the Optimal Sum of Losses
and that Proposition 2.1 yields (8). The next theorem identifies its exact logarithmic asymptotic rate with the weighted Chernoff information from Definition 2.4.
@@
Consider the binary hypotheses $H_{0}:\ X_{1}^{n}\sim\mathbb{P}^{\otimes n}$ versus $H_{1}:\ X_{1}^{n}\sim\mathbb{Q}^{\otimes n}$, under Assumption 1.1. Assume also
$\sup\limits_{\alpha\in[0,1]}\int\varphi(x)|\ln\frac{p(x)}{q(x)}|p(x)^{\alpha}q(x)^{1-\alpha}{\mathrm{d}}\mu(x)<\infty.$ (14)
Let $\mathcal{D}$ be the class of (possibly randomised) decision rules $D:{\cal X}^{n}\to[0,1]$, and let $\alpha_{\varphi}(D)$, $\beta_{\varphi}(D)$ be defined by (6)–(7). Assume that $p,q>0$ $\mu$-a.e. and that $\rho^{\mathrm{w}}_{\alpha}(p,q)\in(0,\infty)$ for all $\alpha\in[0,1]$.
@@
$L_{n}^{*}=\exp\{-nD_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})+o(n)\}.$ (15)
@@
For any $\alpha\in[0,1]$, $\min\{a,b\}\leq a^{\alpha}b^{1-\alpha}$, hence, by factorisation,
@@
Taking the infimum over $\alpha\in[0,1]$ gives
$\liminf_{n\to\infty}-\frac{1}{n}\ln L_{n}^{*}\geq D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q}).$
Now fix $\alpha^{*}\in\arg\min_{\alpha\in[0,1]}\rho^{\mathrm{w}}_{\alpha}(p,q)$ and define
$r_{\alpha^{*}}(x):=\frac{\varphi(x)p(x)^{\alpha^{*}}q(x)^{1-\alpha^{*}}}{\rho^{\mathrm{w}}_{\alpha^{*}}(p,q)},\qquad S_{n}:=\sum_{i=1}^{n}\ln\frac{p(X_{i})}{q(X_{i})}.$
A direct change of measure yields
$L_{n}^{*}=\bigl(\rho^{\mathrm{w}}_{\alpha^{*}}(p,q)\bigr)^{n}\mathbb{E}_{r_{\alpha^{*}}^{\otimes n}}\!\Big[e^{-\alpha^{*}S_{n}}\mathbf{1}\{S_{n}>0\}+e^{(1-\alpha^{*})S_{n}}\mathbf{1}\{S_{n}\leq 0\}\Big].$
The bracket is bounded above by $1$.
Let $F(\alpha):=\ln\rho^{\mathrm{w}}_{\alpha}(p,q)$, it is easy to check that $F(\alpha)$ is a convex function. Under the regularity assumption of (14),
$F^{\prime}(\alpha)=\mathbb{E}_{r_{\alpha}}\!\left[\ln\frac{p(X)}{q(X)}\right].$
If $\alpha^{*}\in(0,1)$, then $F^{\prime}(\alpha^{*})=0$. Hence, in view of LLN, $S_{n}/n\to 0$ in $r_{\alpha^{*}}$-probability. Therefore, for every $\varepsilon>0$,
$\mathbb{E}_{r_{\alpha^{*}}^{\otimes n}}\!\Big[e^{(1-\alpha^{*})S_{n}}\mathbf{1}\{S_{n}\leq 0\}+e^{-\alpha^{*}S_{n}}\mathbf{1}\{S_{n}>0\}\Big]\geq e^{-\varepsilon n}\,r_{\alpha^{*}}^{\otimes n}\big(|S_{n}|\leq\varepsilon n\big)=\exp\{o(n)\}.$
Combined with the upper bound by $1$, this implies
$L_{n}^{*}=\exp\{-nD_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})+o(n)\}.$
In the boundary case $\alpha^{*}=0$, we have the mean value $m>0$, and $\mathbf{1}(S_{n}>0)e^{-\alpha^{*}S_{n}}|_{\alpha^{*}=0}\to 1$ a.s. as $n\to\infty$ by the strong LLN. Similarly, in the case $\alpha^{*}=1$, we have the mean value $m\leq 0$, and $\mathbf{1}(S_{n}\leq 0)e^{(1-\alpha^{*})S_{n}}|_{\alpha^{*}=1}\to 1$ a.s. as $n\to\infty$. This completes the proof. ∎
@@
${\rm TV}_{\varphi}(\mathbb{P}^{\otimes n},\mathbb{Q}^{\otimes n})=\frac{1}{2}\Big((E_{\varphi}(p))^{n}+(E_{\varphi}(q))^{n}\Big)-\exp\{-nD^{\mathrm{w}}_{C}(\mathbb{P},\mathbb{Q})+o(n)\},\penalty 10000\ \penalty 10000\ n\to\infty,$ (16)
@@
Combine the identity (10) with (15). ∎
@@
The weighted $\alpha$-skewed Bhattacharyya distance (12) appears in many papers, see e.g., [12]. Definition 2.4 shows that the weighted Chernoff information divergence is the maximally skewed weighted Bhattacharyya distance.
@@ 3.2 Exponential-Family Representation and Uniqueness of $\alpha^{*}$
$Z_{pq}(\alpha):=\int_{\cal X}\varphi(x)\,p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}\mu(x)=\rho^{\mathrm{w}}_{\alpha}(p,q),$ (17)
@@
${\cal E}_{pq}=\left\{(pq)_{\alpha}(x):=\frac{\varphi(x)\,p(x)^{\alpha}q(x)^{1-\alpha}}{Z_{pq}(\alpha)}\;:\;\alpha\in[0,1]\right\}.$ (18)
@@
Set $t(x):=\ln\frac{p(x)}{q(x)}$ and $k_{pq}(x):=\ln\varphi(x)+\ln q(x)$. Then, $(pq)_{\alpha}$ admits the exponential-family form
$(pq)_{\alpha}(x)$ $=\exp\!\left\{\alpha\,t(x)-F_{pq}(\alpha)+k_{pq}(x)\right\},$ (19)
$F_{pq}(\alpha):$ $=\ln Z_{pq}(\alpha)=-D^{\mathrm{w}}_{B,\alpha}(p,q).$ (20)
@@
The log-normaliser $F_{pq}$ is convex on $[0,1]$; if $\ln\frac{p}{q}$ is not $\mu$-a.e. constant on $\{\varphi>0\}$, then $F_{pq}$ is strictly convex and the maximiser $\alpha^{*}$ in (13) is unique. By Hölder’s inequality,
@@
Finally, note that $Z_{pq}(1)=E_{\varphi}(p)$ and $Z_{pq}(0)=E_{\varphi}(q)$; hence,
@@
so ${\cal E}_{pq}$ is an exponential arc between the tilted versions of $\mathbb{P}$ and $\mathbb{Q}$. By this definition, the following identities hold:
$\begin{array}[]{l}D^{\mathrm{w}}_{C}(p,q)=D^{\mathrm{w}}_{B,\alpha^{*}(p,q)}(p,q)=D^{\mathrm{w}}_{B,\alpha^{*}(q,p)}(q,p)=D^{\mathrm{w}}_{C}(q,p).\end{array}$ (21)
@@ 3.3 Weighted Bregman Divergence and Information-Geometric Identities
3.3.1 Weighted KL Divergence and Weighted Bregman Divergence
@@
$p_{\theta}(x)=\exp\{\theta^{\mathrm{T}}t(x)-F(\theta)+k(x)\},\qquad x\in{\cal X}.$ (22)
For a density $r$, set $E_{\varphi}(r):=\int_{\cal X}\varphi(x)\,r(x)\,\mathrm{d}\mu(x)$ and write $E_{\varphi}(\theta):=E_{\varphi}(p_{\theta})$.
@@
$\hat{F}(\theta):=\ln\int_{\cal X}\varphi(x)\,e^{\theta^{\mathrm{T}}t(x)+k(x)}\,\mathrm{d}\mu(x)=F(\theta)+\ln E_{\varphi}(\theta).$ (23)
@@
$p_{\theta}^{*}(x)=\frac{\varphi(x)\,p_{\theta}(x)}{E_{\varphi}(\theta)}.$ (24)
@@
$D^{\mathrm{w}}_{\rm KL}(p\|q):=\int_{\cal X}\varphi(x)\,p(x)\,\ln\frac{p(x)}{q(x)}\,\mathrm{d}\mu(x),$ (25)
@@
$B^{\mathrm{w}}_{\varphi,F}(\theta_{1},\theta_{2}):$ $=e^{\hat{F}(\theta_{2})-F(\theta_{2})}\Big[F(\theta_{1})-F(\theta_{2})-(\theta_{1}-\theta_{2})^{\mathrm{T}}\nabla\hat{F}(\theta_{2})\Big]$ (26)
@@
For a regular exponential family ${\cal E}=\{\mathbb{P}_{\theta},\theta\in\Theta\}$, assume that the integral in (25) is well-defined in $(-\infty,\infty]$. Then for any $\theta_{1},\theta_{2}\in\Theta$,
$D^{\mathrm{w}}_{\rm KL}(p_{\theta_{1}}\|p_{\theta_{2}})=B^{\mathrm{w}}_{\varphi,F}(\theta_{2},\theta_{1}).$ (27)
@@
This identity is stated in [3, Proposition 4.1]; we give a short derivation for completeness. By (22),
@@
Substituting this into (25) yields
@@
Hence, $\int\varphi tp_{\theta_{1}}\,\mathrm{d}\mu=E_{\varphi}(\theta_{1})\nabla\hat{F}(\theta_{1})$, and therefore
@@
(a) Weighted one-parameter identity. Assume $d=1$ (one-parameter case) and let $\theta_{i}^{*}:=F^{\prime}(\theta_{i})$. Then, the following weighted analogue of the classical Bregman duality holds:
$B^{\mathrm{w}}_{\varphi,F}(\theta_{1},\theta_{2})=B^{\mathrm{w}}_{\varphi,F^{*}}(\theta_{2}^{*},\theta_{1}^{*})-(\theta_{1}-\theta_{2})\,\ln E_{\varphi}(\theta_{2})+(\theta_{2}^{*}-\theta_{1}^{*})\,\ln E_{\varphi}(\theta_{1}),$ (28)
@@
$B_{F}(\theta_{0},\theta_{1})=F(\theta_{0})+F^{*}(\theta_{1}^{*})-\theta_{0}^{\mathrm{T}}\theta_{1}^{*},\qquad\theta_{1}^{*}=\nabla F(\theta_{1}),$ (29)
@@
Part (a) is a weighted extension of the classical duality $B_{F}(\theta_{1},\theta_{2})=B_{F^{*}}(\theta_{2}^{*},\theta_{1}^{*})$ and follows by combining the weighted representation (26) with Legendre relations; see also [3]. Part (b) is standard. ∎
3.3.2 Weighted Chernoff/Bhattacharyya Quantities Inside an Exponential Family